OpenStudy (hesan):

A father is twice as old as his son, six years ago father was thrice as old as his son. What is the age of the son?

OpenStudy (anonymous):

f=2s f+6=3s. -> f=3s-6 3s-6=2s s=6

OpenStudy (ganpat):

F = 2S.. 6 years ago , F = 3( S - 6) therefore 2S = 3S - 18.. S = 18 years.. @student92 : shoud`t be it like this ??

OpenStudy (anonymous):

oh sorry i though he write six years later

OpenStudy (hesan):

But the answer is 6 years...

OpenStudy (hesan):

?

OpenStudy (ganpat):

6 is right , u mean to say ?

OpenStudy (hesan):

yeah

OpenStudy (ganpat):

@mathslover , @hero, @ParthKohli : Guys little confusion here.. can u please help ??

mathslover (mathslover):

A father is twice as old as his son, six years ago father was thrice as old as his son. What is the age of the son? this is the question : Let son's age be x and father's will be : 2x six years ago there age was : x-6 and 2x-6 and given (2x-6)=3(x-6) 2x-6=3x-18 2x-3x=-12 x=12

mathslover (mathslover):

2x = 24

mathslover (mathslover):

father's age = 24 son's age =12 u can verify this also

OpenStudy (hesan):

alright its 12 then..

OpenStudy (ganpat):

ohkey, i missed to subtract father`s age by 6 !! thanks @mathslover ; yes brother.. this is the answer..